# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5.7.2 Datasets
Table 28: List of public datasets in the ”Generic” papers and their corresponding links.
@@
Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90] https://github.com/laiguokun/multivariate-time-series-data
@@
Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85] https://github.com/zhouhaoyi/ETDataset
Energy [art_77] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
@@
Eu-Core [art_202] https://doi.org/10.1145/ (source paper)
Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90] https://github.com/laiguokun/multivariate-time-series-data
@@
METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160] https://paperswithcode.com/dataset/metr-la
@@
NOAA Weather [art_82] https://www.ncei.noaa.gov/data/local-climatological-data/
@@
PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88] https://paperswithcode.com/dataset/pems04
PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_72], [art_75], [art_160] https://paperswithcode.com/dataset/pems-bay
@@
Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90] https://github.com/laiguokun/multivariate-time-series-data
@@
Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90] https://github.com/laiguokun/multivariate-time-series-data
@@
It should be noted that there are multiple recurring datasets that can be considered as benchmarks for the community. Some of these have already been discussed in the previous sections. If the results for these datasets have already been shown in previous sections, they will not be repeated again here. Instead, they will be just marked with a star together with the benchmark models in the tables of the previous sections.
@@
The four most common datasets are Electricity consumption, Exchange-rate, Solar-Energy and Traffic, all available at https://github.com/laiguokun/multivariate-time-series-data. The Electricity consumption dataset contains the hourly electricity consumption in kWh for 321 customers recorded from 2012 to 2014. The Traffic dataset describes the road occupancy rates measured by different sensors on the San Francisco Bay area road network from 2015 to 2016, with an hourly granularity. In Solar energy it is recorded the solar power production from 137 photovoltaic plants in Alabama State in 2006, with a 10 minutes granularity. Finally, Exchange rate contains daily exchange rates of 8 countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore, from 1990 to 2016.
@@
In general, the data granularity varies significantly, ranging from fractions of seconds to one week, indicating that there is no particular focus on a specific granularity. Individual papers often cover multiple granularities, and many papers take into account datasets with granularity from 5 minutes to 1 day. As for the forecasting horizon, almost all papers focus on multi-step forecasting, usually considering 12 different forecasting horizons at the same time. Only a few papers mention data pre-processing, and usually utilize min-max or Z-score normalization.
5.7.3 Proposed models
@@
Among the 24 ”Generic” papers, 19 use a purely convolutional GNN approach, and two employ a purely attentional model. One paper lacks sufficient detail to determine its classification within the taxonomy [art_85], and another declares it is using an aggregating approach that is simpler than either convolutional or attentional approaches [art_81]. One more paper proposes a multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75].
@@
As for the definition of the graph structure, the majority of these models use self-learning techniques. This is a straightforward choice, since all the papers in this section want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability. Half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
@@
The two most commonly utilized loss functions are the MSE and the MAE. Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs. [art_73], [art_78], [art_79], [art_86]).
@@
Table 29: List of source codes of the ”Generic” models in the review.
@@ 5.7.4 Benchmark models
Tab. LABEL:t:benchmark_models_generic displays the benchmark models utilized by at least two of the selected ”Generic” papers.
Table 30: List of benchmark models in the ”Generic” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89]
@@
Historical average (HA) [art_78], [art_88]
Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88]
Classical machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78]
@@
Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90]
@@
Informer [Informer] [art_74], [art_76], [art_82], [art_85]
@@
Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89]
Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90]
@@
Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87]
@@
Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90]
Transformer [attention] [art_71], [art_77]
Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88]
@@
Graph for time series (GTS) [GTS] [art_72], [art_78]
Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90]
@@
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78]
There is a wide variety of benchmark models in this ”Generic” group. As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability. Among classical machine learning models, the most widely used models are LSTMs, temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and long-short term time series network (LSTNet) [LSTNet]. The TPA-LSTM model, introduced by Shih et al. in 2019, employs a set of filters to capture time-invariant temporal patterns, and an attention mechanism to identify relevant time series for multivariate forecasting. The source code for this benchmark model is available at https://github.com/gantheory/TPA-LSTM. The LSTNet model, proposed by Lai et al. in 2018, combines a convolution neural network and a recurrent neural network to capture short-term local dependencies among variables and identify long-term patterns. Additionally, it incorporates an autoregressive model that enhances the robustness of the deep learning approach to time series with significant scale fluctuations. The source code for the model is accessible at https://github.com/laiguokun/LSTNet. Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN], already discussed in the previous sections.

### v2

@@ 5.7.2 Datasets
Table 33: List of public datasets in the “Generic” papers and their corresponding links.
@@
CDC US States [art_401] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
@@
C-MAPSS [art_403] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
CORAv2.0 [art_396] https://mds.nmdis.org.cn/pages/home.html
@@
ECG5000 [art_397] https://www.cs.ucr.edu/ẽamonn/time_series_data/
Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90], [art_384], [art_405] https://github.com/laiguokun/multivariate-time-series-data
@@
Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85], [art_393], [art_417] https://github.com/zhouhaoyi/ETDataset
Energy [art_77], [art_388] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
@@
Eu-Core [art_202] https://doi.org/10.1145/3018661.3018731 (source paper)
Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90], [art_384], [art_393], [art_405] https://github.com/laiguokun/multivariate-time-series-data
@@
IDWR Japan Prefectures [art_401] https://www.niid.go.jp/niid/en/idwr-e.html
@@
ISRUC-SLEEP Dataset [art_403] https://sleeptight.isr.uc.pt/ISRUC_Sleep/
@@
KnowAir [art_400] https://github.com/shuowang-ai/PM2.5-GNN
LA Traffic [art_404] https://ladot.lacity.gov/residents/traffic-counts
@@
Los-loop [art_383] https://github.com/lehaifeng/T-GCN/tree/master/data
Large-Scale Traffic and Weather Events (LSTW) [art_404] https://smoosavi.org/datasets/lstw
METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160], [art_387], [art_397], [art_401], [art_417], [art_464] https://paperswithcode.com/dataset/metr-la
Metro [art_400] https://lbs.amap.com/
@@
NOAA Weather [art_82], [art_393] https://www.ncei.noaa.gov/data/local-climatological-data/
NYC-Taxi [art_398] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
@@
PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems04
PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_72], [art_75], [art_160], [art_401] https://paperswithcode.com/dataset/pems-bay
@@
Seattle-loop [art_401] https://github.com/zhiyongc/Seattle-Loop-Data
Sleep-EDF-20 [art_395] https://gist.github.com/emadeldeen24/a22691e36759934e53984289a94cb09b
Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90], [art_384], [art_387], [art_405] https://github.com/laiguokun/multivariate-time-series-data
Secure Water Treatment (SWaT) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
Stocktwits [art_398] https://stocktwits.com/
@@
Taiwan PM2.5 [art_464] https://taqm.epa.gov.tw/taqm/en/default.aspx
T-drive [art_383] https://www.kaggle.com/datasets/arashnic/tdriver
Tiingo [art_398] https://www.tiingo.com/
Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90], [art_388] https://github.com/laiguokun/multivariate-time-series-data
UCI Human Activity Recognition [art_403] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones
University of East Anglia (UEA) datasets [art_385], [art_386], [art_389], [art_407] https://timeseriesclassification.com/
@@
Water Distribution (WaDi) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
WESAD dataset [art_404] https://ubi29.informatik.uni-siegen.de/usi/data_wesad.html
@@
Wind Generation [art_405] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
WISDM [art_395] https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset)
@@
It should be noted that multiple recurring datasets can be considered as benchmarks for this community. Some of these were discussed in the previous subsections (e.g., METR-LA, PeMS datasets). Hence, if the results for these datasets were already shown in previous subsections, they will not be discussed again here. Instead, they will be just marked with the symbol in the tables of the previous subsections.
@@
The four most common datasets are Electricity consumption, Exchange-rate, Solar-Energy and Traffic, all available at https://github.com/laiguokun/multivariate-time-series-data. The Electricity consumption dataset contains the hourly electricity consumption in kWh for 321 customers, recorded from 2012 to 2014. The Traffic dataset describes the road occupancy rates measured by different sensors on the San Francisco Bay area road network from 2015 to 2016, with an hourly granularity. In Solar energy it is recorded the solar power production from 137 photovoltaic plants in Alabama State in 2006, with a 10 minutes granularity. Finally, Exchange rate contains daily exchange rates of 8 countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore, from 1990 to 2016.
The University of East Anglia (UEA) dataset includes 30 multivariate time series from many real-world applications, including heart records, electroencephalogram signals, motion data, audio spectra, written characters. Since a large portion of the dataset contains data not related to the focus of this review, the details of the results and models used will not be discussed here. Interested readers are referred to Refs. [art_385, art_386, art_389, art_407].
@@
In general, the data granularity varies significantly, ranging from fractions of seconds to one month, indicating that there is no particular focus on a specific granularity. Individual papers often cover multiple granularities, and many papers take into account datasets with granularity from 5 minutes to 1 day. As for the forecasting horizon, almost all papers focus on multi-step forecasting, usually considering 12 different forecasting horizons at the same time. Only a few papers mention data pre-processing, and usually utilize min-max or Z-score normalization.
5.7.3 Types of models
@@
Among the 43 papers in the “Generic” thematic group, 33 use a purely convolutional GNN approach, and four employ a purely attentional model. Five paper lack sufficient detail to be classified within the taxonomy chosen (Refs. [art_85, art_385, art_386, art_389, art_407]), and another one declares that it uses an aggregating approach that is simpler than either convolutional or attentional approaches [art_81]. Another paper proposes a so-called multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75]. This is not the only paper in this section using differential equations. In Ref. [art_401], two domain specific ODEs (one for traffic speed evolution, and the other for disease propagation) are integrated into the convolutional GNN model. The authors argue that this dynamics-informed model approach enhances robustness, allow easier generalization when the equation of the dynamics is known, and reduces the number of model parameters.Additionally, during inference, the model only requires a single data point, as the ODEs explicitly incorporate the immediate dynamics.
@@
As for the definition of the graph structure, the majority of these models use self-learning techniques. This is a straightforward choice, since all the papers in this group want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability. More than half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
@@
The two most commonly utilized loss functions are the MSE and the MAE. Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs. [art_73, art_78, art_79, art_86]).
@@
Table 34: List of source codes of the “Generic” models in the review.
@@
Attention-adjusted Graph Spatio-Temporal Network (AGSTN) [art_464] https://github.com/l852888/AGSTN
Cross-attention temporal dynamic graph neural network (CATodyNet) [art_386] https://github.com/HQYWY/CATodyNet
@@
Dynamic diffusion-variational GNN (DVGNN) [art_383] https://github.com/gorgen2020/DVGNN
@@
GraphSensor [art_395] https://github.com/jldx-gjq/Graphnet/tree/main/graphsensor-code-main
@@
Multi-task Dynamic Graph generation framework for multivariate time series forecasting (MDG) [art_397] https://github.com/gortwwh/MDG
@@
Multivariate Time Series Classification based on Fusion Features (MSTC_FF) [art_385] https://github.com/dumingsen/MTSC_FF
@@
Temporal Dynamic Graph Neural Network (TodyNet) [art_389] https://github.com/liuxz1011/TodyNet
@@ 5.7.4 Benchmark models
Tab. LABEL:t:benchmark_models_generic displays the benchmark models utilized by at least two of the selected “Generic” papers.
Table 35: List of benchmark models in the “Generic” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89], [art_384], [art_393], [art_397], [art_400], [art_405], [art_464]
@@
Historical average (HA) [art_78], [art_88], [art_400]
Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88], [art_384], [art_400], [art_405]
Non-GNN machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
DeepAR [DeepAR] [art_69], [art_393]
DeepSleepNet [DeepSleepNet] [art_395], [art_403]
Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78], [art_405]
@@
Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90], [art_384], [art_398], [art_404], [art_405]
@@
Informer [Informer] [art_74], [art_76], [art_82], [art_85], [art_397]
$k$-nearest neighbors (KNN) [art_84], [art_387]
@@
Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89], [art_396], [art_397], [art_398], [art_400], [art_464]
Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90], [art_387], [art_393]
@@
Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_405]
@@
Support vector machine (SVM) [art_84], [art_396], [art_464]
Temporal convolutional network (TCN) [TCN] [art_88], [art_387], [art_396], [art_397]
Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90], [art_405]
Transformer [attention] [art_71], [art_77], [art_393]
Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90], [art_405]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88], [art_384], [art_396], [art_400], [art_464]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79], [art_383], [art_384], [art_396], [art_397], [art_400], [art_464]
Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_74], [art_396]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405]
Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] [art_397], [art_398]
Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_396], [art_417]
Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) [DSTAGNN] [art_383], [art_384]
@@
Discrete graph structure learning for time series (GTS) [GTS] [art_72], [art_78], [art_384], [art_397], [art_401]
Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87], [art_396], [art_397], [art_417]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90], [art_393], [art_396], [art_397], [art_401], [art_405], [art_417]
Multivariate Time series with dynamic Graph neural Ordinary Differential Equations (MTGODE) [art_75] [art_88], [art_401], [art_417]
@@
Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_397], [art_405]
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86], [art_405]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88], [art_384], [art_397], [art_400], [art_401]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88], [art_383]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78], [art_387], [art_397], [art_400], [art_405]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_383], [art_384], [art_393]
Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_400], [art_405]
This “Generic” thematic group includes a wide variety of benchmark models, also due to the large number of papers within this group. As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability. Among non-GNN machine learning models, the most widely used models are LSTM networks, the so-called temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and the long-short term time series network (LSTNet) [LSTNet]. The TPA-LSTM model, introduced by Shih et al. in 2019, employs a set of filters to capture time-invariant temporal patterns, and an attention mechanism to identify relevant time series for multivariate forecasting. The source code for this benchmark model is available at https://github.com/gantheory/TPA-LSTM. The LSTNet model, proposed by Lai et al. in 2018, combines a convolution neural network and a recurrent neural network to capture short-term local dependencies among variables and identify long-term patterns. Additionally, it incorporates an autoregressive model that enhances the robustness of the deep learning approach to time series with significant scale fluctuations. The source code for the model is accessible at https://github.com/laiguokun/LSTNet. Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN] models, already discussed in the previous subsections.